# Spin Bott Indices For Time-Reversal-Invariant Higher-Order Topological Superconductors

### v1

@@ IV Edge theory
${H}(\bm{k})=f_{x}(\bm{k})\tau_{z}\sigma_{x}s_{0}+f_{y}(\bm{k})\tau_{z}\sigma_{y}s_{0}+f_{z}(\bm{k})\tau_{z}\sigma_{z}s_{z}+\Delta_{\text{A}}\tau_{x}(\sigma_{0}+\sigma_{z})/2s_{0}+\Delta_{\text{B}}\tau_{x}(\sigma_{0}-\sigma_{z})/2s_{0},$
@@
$f_{z}(\bm{k})=\lambda_{\text{so}}(2\sin(k_{x})-4\sin(k_{x}/2)\cos(\sqrt{3}k_{y}/2)).$ (S59)
@@
${H}_{1}=\Delta_{\text{A}}\tau_{x}(\sigma_{0}+\sigma_{z})/2s_{0}+\Delta_{\text{B}}\tau_{x}(\sigma_{0}-\sigma_{z})/2s_{0}+\sqrt{3}tk_{y}\cos(k_{x}/2)\tau_{z}\sigma_{y}s_{0}.$ (S60)
@@
$\psi_{4}=|\tau_{z}=1\rangle\otimes|\sigma_{y}=-1\rangle\otimes|s_{z}=-1\rangle.$ (S61)
The $k_{y}$-dependent term couples the zero-energy states away from $k_{y}=0$, which gives rise to the dispersion of edge state. At exact Dirac point $k_{y}=0$, the above four zero-energy states are coupled by the term $\Delta_{\text{A}}/2\tau_{x}\sigma_{0}s_{0}+\Delta_{\text{B}}/2\tau_{x}\sigma_{0}s_{0}$, which leads to the superconducting gap with magnitude $(\Delta_{\text{A}}+\Delta_{\text{B}})/2$. This is consistent with the numerical results shown in Fig. S1 (a) and S1(b).
@@
${H}_{1}=-2\lambda_{\text{so}}\delta k_{x}(1-\cos(\sqrt{3}k_{y}/2))\tau_{z}\sigma_{z}s_{z}+\Delta_{\text{A}}\tau_{x}(\sigma_{0}+\sigma_{z})/2s_{0}+\Delta_{\text{B}}\tau_{x}(\sigma_{0}-\sigma_{z})/2s_{0}.$ (S62)
@@
Here $\delta k_{x}=k_{x}-\pi$. As ${H}_{0}$ only contains $\sigma_{x}$ and $\sigma_{y}$ in the sublattice space, the zero-energy states of ${H}_{0}$ are the eigenstates of $|\sigma_{z}=1\rangle$ ($|\sigma_{z}=-1\rangle$) when the zigzag edge is formed by the A (B) atom. The $\lambda_{\text{so}}$-dependent term couples the zero-energy states away from $k_{x}=\pi$, which gives rise to the dispersion of edge states. However, at exact Dirac point $k_{x}=\pi$, $\lambda_{\text{so}}$-dependent term vanishes and the zero-energy states are coupled by the superconducting pairing, leading to the superconducting gap with magnitude $\Delta_{\text{A}}$ ($\Delta_{\text{B}}$) when the zigzag edge is formed by the A (B) atom. For the bearded edge, the boundary Dirac points are located at $k_{x}=0$, which are also gapped by the superconducting pairing with magnitude $\Delta_{\text{A}}$ ($\Delta_{\text{B}}$) when the bearded edge is formed by the A (B) atom. The zigzag and bearded edge states spectra are shown in Figs. S1 (c-d) and Figs. S1 (e-f), respectively.
@@
V Majoaran Kramers Pairs in the square- and dodecagonal-shaped systems
@@
As explained in the main text, Majoaran Kramers Pairs (MKP) can be realized at a corner shared by armchair and zigzag (bearded) edges when $(\Delta_{\text{A}}+\Delta_{\text{B}})\Delta_{\alpha}<0$ (the zigzag (bearded) edge is formed by the $\alpha$ atom). As the angles between the armchair and zigzag (bearded) edges can take the values of $\pi/2$ and $65\pi/$ 6, we can engineer MKP by tailoring the systems into square and dodecagonal shapes. For square and dodecagonal systems, the edge-state energy gaps of armchair edges have the same sign, which limits possible patterns of MKP realized through edge cleavage in these systems.
Figure S2: Different patterns of MKP on the square-shaped systems. The circled number in (a) at corners labels the order of corners. The letters $\text{A}_{\text{AB}}$, $\text{Z}_{\alpha}$, $\text{B}_{\alpha}$ label the armchair, zigzag, and bearded edges, respectively, with $\alpha\in$A, B}. The insets plot the energies close to zero. The vectors $\bm{\chi}_{4}$ and $\bm{\mathcal{N}}_{4}$ are used to describe and characterize the different patterns of MKP. The model parameters are taken as $t=1,\lambda_{so}=0.2,\Delta_{\text{A}}=0.5,\Delta_{\text{B}}=-0.25$.
@@
For a square-shaped system, MKP can be realized at two adjacent corners [Fig. S2(a)] or all four corners [Fig. S2(b)]. To characterize the two patterns of MKP described by the $\bm{\chi}_{4}$ vector [Fig. S2], we calculate the spin Bott index ${\mathcal{N}}_{4}^{(1,2,3)}$, which are generated by choosing
${f}_{4}^{(1)}=2xy/L^{2},{f}_{4}^{(2)}=x/L,{f}_{4}^{(3)}=y/L,$ (S63)
@@
where $(x,y)$ denotes the coordinate of lattice sites and $L$ is the side length. In Figs. S2 (a) and S2(b), we present the numerical values of ${\mathcal{N}}_{4}$ which fully capture the different patterns of MKP described by $\bm{\chi}_{4}=\mathcal{M}^{-1}\cdot\left({\mathcal{N}}_{4}^{(1)},{\mathcal{N}}_{4}^{(2)},{\mathcal{N}}_{4}^{(3)},0\right)^{\mathrm{T}}$ with
$\mathcal{M}=\frac{1}{2}\begin{pmatrix}1&-1&1&-1\\ 1&-1&-1&1\\ 1&1&-1&-1\\ 1&1&1&1\end{pmatrix}.$ (S64)
Figure S3: (a-f) Different patterns of MKP on dodecagonal-shaped systems. The letters $\text{A}_{\text{AB}}$, $\text{Z}_{\alpha}$, $\text{B}_{\alpha}$ label the armchair, zigzag, and bearded edges, respectively, with $\alpha\in$A, B}. The circled number in (a) at corners labels the order of corners. The insets plot the energies close to zero. The vectors $\bm{\chi}_{12}$ and $\bm{\mathcal{N}}_{12}$ are used to describe and characterize the different patterns of MKP. The same model parameters are used as those in Fig. S2.
@@
For the dodecagon-shaped system, MKP can be realized at $q$ corners by edge cleavage with $q=2,4,6,8,10,12$, as exemplified in Figs. S3(a)- S3(f). According to our theory, eleven spin Bott indices ${\mathcal{N}}_{12}^{(1,\cdots,11)}$ are needed to fully characterize the different patterns of MKP. The invariants ${\mathcal{N}}_{12}^{(1,\cdots,11)}$ can be generated by choosing [60]
@@
${f}_{12}^{(11)}=-(1/3)\sqrt{2}(x-\sqrt{3}y)(x^{4}+7\sqrt{3}x^{3}y+7x^{2}y^{2}-\sqrt{3}xy^{3}-2y^{4})/2L^{5},$ (S65)
where $L$ is the length between the origin and a corner. In Figs. S3(a)-S3(f), we present the numerical values of ${\mathcal{N}}_{12}$ which fully characterize the different patterns of MKP described by $\bm{\chi}_{12}$ and we have $\bm{\chi}_{12}=\mathcal{M}^{-1}\cdot\left({\mathcal{N}}_{12}^{(1)},\cdots,{\mathcal{N}}_{12}^{(11)},0\right)^{\mathrm{T}}$ with
$\mathcal{M}=\frac{1}{2}\left(\begin{array}[]{cccccccccccc}-1&1&-1&1&-1&1&-1&1&-1&1&-1&1\\ 1&1&1&-1&-1&-1&1&1&1&-1&-1&-1\\ -1&-1&1&1&-1&1&-1&-1&1&1&-1&1\\ -1&-1&1&1&1&-1&-1&-1&1&1&1&-1\\ 1&-1&1&1&-1&-1&1&-1&1&1&-1&-1\\ -1&-1&-1&-1&-1&-1&1&1&1&1&1&1\\ 1&1&1&1&1&-1&-1&-1&-1&-1&-1&1\\ 1&1&1&1&-1&1&-1&-1&-1&-1&1&-1\\ 1&1&1&-1&1&1&-1&-1&-1&1&-1&-1\\ -1&-1&1&-1&-1&-1&1&1&-1&1&1&1\\ -1&1&-1&-1&-1&-1&1&-1&1&1&1&1\\ 1&1&1&1&1&1&1&1&1&1&1&1\\ \end{array}\right).$ (S66)
Figure S4: Systems in diamond (a) and honeycomb shapes (b) without MKP. The insets plot the energies close to zero. The same model parameters are used as those in Fig. 1.
VI Absence of Majorana Kramers Pair
In the diamond- and honeycomb-shaped systems, every edge belongs to the type-I edge (armchair or bearded edge) and is formed by A or B atoms. When all the edges are formed by the same atom, there are no MKP of corner states, as shown in Fig. S4.
In the square- and dodecagon-shaped systems, adjacent edges belong to different types of edges and are gapped with the magnitude $(\Delta_{\text{A}}+\Delta_{\text{B}})/2$ and $\Delta_{\alpha}$, respectively. By tuning $\Delta_{\text{A}}$ and $\Delta_{\text{B}}$ to satisfy $(\Delta_{\text{A}}+\Delta_{\text{B}})\Delta_{\alpha}>0$, there are no MKP.
Here we want to emphasize that the corner states are sensitive to the edge termination.
VII $Z_{4}$ symmetry indicator
$\kappa=\frac{1}{4}\sum_{\bm{k}\in\mathrm{TRIM}}\sum_{n}\xi_{\bm{k},n}\,\,\,\,\text{mod}\,\,4,$ (S67)
where $\xi_{\bm{k},n}$ is the parity eigenvalue of the occupied $n$th BdG band at time-reversal invariant momenta $\bm{k}$ and $\kappa=2$ characterizes the inversion-protected MKP [37]. In Table S1, we present the parity eigenvalue of each occupied BdG band at the four time-reversal invariant momenta and we have $\kappa=2$ by applying Eq. (S67) to ${H}(\bm{k})$.
Table S1: The inversion eigenvalues $\xi_{\bm{k},n}$ of all the four occupied BdG bands $n=1,2,3,4$, at $\bm{k}=\bm{\Gamma},\bm{M_{1}},\bm{M_{2}},\bm{M_{3}}$, with $\bm{\Gamma}=(0,0),\bm{M_{1}}=(0,\frac{2\pi}{\sqrt{3}}),\bm{M_{2}}=(\pi,\frac{\pi}{\sqrt{3}})$ and $\bm{M_{3}}=(-\pi,\frac{\pi}{\sqrt{3}})$.

### v3

@@ Appendix E Edge theory
${H}(\bm{k})=f_{x}(\bm{k})\tau_{z}\sigma_{x}s_{0}+f_{y}(\bm{k})\tau_{z}\sigma_{y}s_{0}+f_{z}(\bm{k})\tau_{z}\sigma_{z}s_{z}+$
$\quad\quad\quad\quad\Delta_{\text{A}}\tau_{x}(\sigma_{0}+\sigma_{z})/2s_{0}+\Delta_{\text{B}}\tau_{x}(\sigma_{0}-\sigma_{z})/2s_{0},$
@@
$f_{z}(\bm{k})=\lambda_{\text{so}}(2\sin(k_{x})-4\sin(k_{x}/2)\cos(\sqrt{3}k_{y}/2)).$ (72)
@@
Figure 7: (a-f) Different patterns of MKP on dodecagonal-shaped systems. The letters $\text{A}_{\text{AB}}$, $\text{Z}_{\alpha}$, $\text{B}_{\alpha}$ label the armchair, zigzag, and bearded edges, respectively, with $\alpha\in$A, B}. The circled number in (a) at corners labels the order of corners. The insets plot the energies close to zero. The vectors $\bm{\chi}_{12}$ and $\bm{\mathcal{N}}_{12}$ are used to describe and characterize the different patterns of MKP. The same model parameters are used as those in Fig. 6.
@@
${H}_{1}=\Delta_{\text{A}}\tau_{x}(\sigma_{0}+\sigma_{z})/2s_{0}+\Delta_{\text{B}}\tau_{x}(\sigma_{0}-\sigma_{z})/2s_{0}+\sqrt{3}tk_{y}\cos(k_{x}/2)\tau_{z}\sigma_{y}s_{0}.$ (73)
@@
$\psi_{4}=|\tau_{z}=1\rangle\otimes|\sigma_{y}=-1\rangle\otimes|s_{z}=-1\rangle.$ (74)
The $k_{y}$-dependent term couples the zero-energy states away from $k_{y}=0$, which gives rise to the dispersion of edge state. At exact Dirac point $k_{y}=0$, the above four zero-energy states are coupled by the term $\Delta_{\text{A}}/2\tau_{x}\sigma_{0}s_{0}+\Delta_{\text{B}}/2\tau_{x}\sigma_{0}s_{0}$, which leads to the superconducting gap with magnitude $(\Delta_{\text{A}}+\Delta_{\text{B}})/2$. This is consistent with the numerical results shown in Fig. 5 (a) and 5(b).
@@
${H}_{1}=-2\lambda_{\text{so}}(\delta k_{x}+2\cos(\delta_{k_{x}})\cos(\sqrt{3}k_{y}/2))\tau_{z}\sigma_{z}s_{z}$
$\quad\quad\quad+\Delta_{\text{A}}\tau_{x}(\sigma_{0}+\sigma_{z})/2s_{0}+\Delta_{\text{B}}\tau_{x}(\sigma_{0}-\sigma_{z})/2s_{0}.$ (75)
@@
Here $\delta k_{x}=k_{x}-\pi$. As ${H}_{0}$ only contains $\sigma_{x}$ and $\sigma_{y}$ in the sublattice space, the zero-energy states of ${H}_{0}$ are the eigenstates of $|\sigma_{z}=1\rangle$ ($|\sigma_{z}=-1\rangle$) when the zigzag edge is formed by the A (B) atom. The term $-4\lambda_{\text{so}}\cos(\delta_{k_{x}})\cos(\sqrt{3}k_{y}/2))\tau_{z}\sigma_{z}s_{z}$ can not remove the zero-energy states but modify the spinor part of the zero-energy states. When $\lambda_{\text{so}}/t$ is small, the zero-energy states approximately take the form $|\sigma_{z}=1\rangle$ or $|\sigma_{z}=-1\rangle$ [35]. For convenience, we do not consider the effect of this term, which does not affect our conclusions. At Dirac point $k_{x}=\pi$, the term $-2\lambda_{\text{so}}\delta k_{x}\tau_{z}\sigma_{z}s_{z}$ vanish and the zero-energy states are coupled by the superconducting pairing, leading to the superconducting gap with magnitude $\Delta_{\text{A}}$ ($\Delta_{\text{B}}$) when the zigzag edge is formed by the A (B) atom. For the bearded edge, the boundary Dirac points are located at $k_{x}=0$, which are also gapped by the superconducting pairing with magnitude $\Delta_{\text{A}}$ ($\Delta_{\text{B}}$) when the bearded edge is formed by the A (B) atom. The zigzag and bearded edge states spectra are shown in Figs. 5 (c-d) and Figs. 5 (e-f), respectively.
@@
Appendix F Majoaran Kramers Pairs in the square- and dodecagonal-shaped systems
@@
As explained in the main text, Majoaran Kramers Pairs (MKP) can be realized at a corner shared by armchair and zigzag (bearded) edges when $(\Delta_{\text{A}}+\Delta_{\text{B}})\Delta_{\alpha}<0$ (the zigzag (bearded) edge is formed by the $\alpha$ atom). As the angles between the armchair and zigzag (bearded) edges can take the values of $\pi/6$ ,$\pi/2$ and $65\pi/$ 6, we can engineer MKP by tailoring the systems into triangular, diamond, square, and dodecagonal shapes. For these systems, the edge-state energy gaps of armchair edges have the same sign, which constrains possible patterns of MKP realized through edge cleavage. In the following, we take the square- and dodecagonal-shaped systems as examples.
@@
For a square-shaped system, MKP can be realized at two adjacent corners [Fig. 6(a)] or all four corners [Fig. 6(b)]. To characterize the two patterns of MKP described by the $\bm{\chi}_{4}$ vector [Fig. 6], we calculate the spin Bott index ${\mathcal{N}}_{4}^{(1,2,3)}$, which are generated by choosing
${f}_{4}^{(1)}=2xy/L^{2},{f}_{4}^{(2)}=x/L,{f}_{4}^{(3)}=y/L,$ (76)
@@
where $(x,y)$ denotes the coordinate of lattice sites and $L$ is the side length. In Figs. 6 (a) and 6(b), we present the numerical values of ${\mathcal{N}}_{4}$ which fully capture the different patterns of MKP described by $\bm{\chi}_{4}=\mathcal{M}^{-1}\cdot\left({\mathcal{N}}_{4}^{(1)},{\mathcal{N}}_{4}^{(2)},{\mathcal{N}}_{4}^{(3)},0\right)^{\mathrm{T}}$ with
$\mathcal{M}=\frac{1}{2}\begin{pmatrix}1&-1&1&-1\\ 1&-1&-1&1\\ 1&1&-1&-1\\ 1&1&1&1\end{pmatrix}.$ (77)
@@
For the dodecagon-shaped system, MKP can be realized at $q$ corners by edge cleavage with $q=2,4,6,8,10,12$, as exemplified in Figs. 7(a)- 7(f). According to our theory, eleven spin Bott indices ${\mathcal{N}}_{12}^{(1,\cdots,11)}$ are needed to fully characterize the different patterns of MKP. The invariants ${\mathcal{N}}_{12}^{(1,\cdots,11)}$ can be generated by choosing [78]
@@
${f}_{12}^{(11)}=-(1/3)\sqrt{2}(x-\sqrt{3}y)(x^{4}+7\sqrt{3}x^{3}y+7x^{2}y^{2}-\sqrt{3}xy^{3}-2y^{4})/2L^{5},$ (78)
where $L$ is the length between the origin and a corner. In Figs. 7(a)-7(f), we present the numerical values of ${\mathcal{N}}_{12}$ which fully characterize the different patterns of MKP described by $\bm{\chi}_{12}$ and we have $\bm{\chi}_{12}=\mathcal{M}^{-1}\cdot\left({\mathcal{N}}_{12}^{(1)},\cdots,{\mathcal{N}}_{12}^{(11)},0\right)^{\mathrm{T}}$ with
$\mathcal{M}=\frac{1}{2}\left(\begin{array}[]{cccccccccccc}-1&1&-1&1&-1&1&-1&1&-1&1&-1&1\\ 1&1&1&-1&-1&-1&1&1&1&-1&-1&-1\\ -1&-1&1&1&-1&1&-1&-1&1&1&-1&1\\ -1&-1&1&1&1&-1&-1&-1&1&1&1&-1\\ 1&-1&1&1&-1&-1&1&-1&1&1&-1&-1\\ -1&-1&-1&-1&-1&-1&1&1&1&1&1&1\\ 1&1&1&1&1&-1&-1&-1&-1&-1&-1&1\\ 1&1&1&1&-1&1&-1&-1&-1&-1&1&-1\\ 1&1&1&-1&1&1&-1&-1&-1&1&-1&-1\\ -1&-1&1&-1&-1&-1&1&1&-1&1&1&1\\ -1&1&-1&-1&-1&-1&1&-1&1&1&1&1\\ 1&1&1&1&1&1&1&1&1&1&1&1\\ \end{array}\right).$ (79)
Table 1: The inversion eigenvalues $\xi_{\bm{k},n}$ of all the four occupied BdG bands $n=1,2,3,4$, at $\bm{k}=\bm{\Gamma},\bm{M_{1}},\bm{M_{2}},\bm{M_{3}}$, with $\bm{\Gamma}=(0,0),\bm{M_{1}}=(0,\frac{2\pi}{\sqrt{3}}),\bm{M_{2}}=(\pi,\frac{\pi}{\sqrt{3}})$ and $\bm{M_{3}}=(-\pi,\frac{\pi}{\sqrt{3}})$.
@@
Appendix G $Z_{4}$ symmetry indicator
$\kappa=\frac{1}{4}\sum_{\bm{k}\in\mathrm{TRIM}}\sum_{n}\xi_{\bm{k},n}\,\,\,\,\text{mod}\,\,4,$ (80)
where $\xi_{\bm{k},n}$ is the parity eigenvalue of the occupied $n$th BdG band at time-reversal invariant momenta $\bm{k}$ and $\kappa=2$ characterizes the inversion-protected MKP [40]. In Table 1, we present the parity eigenvalue of each occupied BdG band at the four time-reversal invariant momenta and we have $\kappa=2$ by applying Eq. (80) to ${H}(\bm{k})$.